OpenStudy (anonymous):

Write 3.75(vinculum above 75) as a reduced fraction

OpenStudy (anonymous):

hi Can you help?

OpenStudy (phi):

The bar over the 75 means it repeats forever

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So if that is, how do you write it as reduced fraction?

OpenStudy (anonymous):

one second, so laggy, stay please

OpenStudy (anonymous):

ok

OpenStudy (valpey):

3.75757575757 = x 375.7575757575 = 100x 372 = 99x 372/99 = x

OpenStudy (phi):

let x= 3.75

OpenStudy (phi):

I see valpey did it.

OpenStudy (anonymous):

Not sure what he means...

OpenStudy (anonymous):

how did it get from 375 to 372

OpenStudy (valpey):

\[= 3 \frac{75}{99}\]

OpenStudy (phi):

x is the original number 100x = 375.75 with 75 repeating

OpenStudy (valpey):

375.75757-3.75757

OpenStudy (anonymous):

ohh

OpenStudy (phi):

when you subtract 100x-x, the repeating 75 cancels

OpenStudy (anonymous):

Wait..

OpenStudy (anonymous):

I remember something about my teacher using x/99, but i dont remember all of it

OpenStudy (valpey):

\[5.123123123123123... = 5 \frac{123}{999}\]

OpenStudy (anonymous):

right, that i understand

OpenStudy (anonymous):

because 99 or 999 will keep it repeating

OpenStudy (anonymous):

But how would it relate to mine?Like how to use that with my question

OpenStudy (anonymous):

What kind of answer is it looking for?

OpenStudy (anonymous):

How would it be 372/99? What exactly is the answer?

OpenStudy (phi):

They would want you to divide it out, to get a mixed number

OpenStudy (phi):

Do you understand how to get to 372/99 ?

OpenStudy (anonymous):

no, sorry i dont understand..

OpenStudy (anonymous):

3.75 with line over it is 3 75/99

OpenStudy (anonymous):

But where does the 372 come from?

OpenStudy (phi):

To get to your answer, you would do what valpey showed. It's a trick to get rid of the repeating part... let x= 3.75 let 100x= 375.75 (in both cases the .75 is repeating) subtract the 2 equations. solve for x

OpenStudy (valpey):

\[375.7575757...-3.757575... = 372\]

OpenStudy (anonymous):

Phi what do you mean subtract the 2 equations? Valpey why is it subtract?

OpenStudy (phi):

100x = 375.7575... - x = - 3.7575.... ------------------ 99 x= 372 (the nasty .7575... goes away)

OpenStudy (anonymous):

Oh, i think i see it

OpenStudy (valpey):

Terrific!

OpenStudy (anonymous):

But it asks as a reduced fraction. Shouldnt it be like a fraction answer?

OpenStudy (valpey):

\[\frac{372}{99} = 3\frac{75}{99}\]

OpenStudy (anonymous):

And so the final answer would be 372/99?

OpenStudy (phi):

I think they would want a mixed number. See http://www.webmath.com/redfract.html

OpenStudy (valpey):

\[3\frac{75}{99} = 3\frac{25}{33}\]

OpenStudy (anonymous):

Oh, because it's more simplified?

OpenStudy (valpey):

Yes, 3 & 25/33 would be the best answer.

OpenStudy (anonymous):

Okay, thank you very much guys, i think i get it

OpenStudy (anonymous):

If you want, can you help me with one quick one?

OpenStudy (valpey):

Just start a new question. It helps other people too.

OpenStudy (anonymous):

ok ok, thanks